Task Time Predecessor 1 0.2 - 2 0.4 1 3 0.1 1 4 0.4 2,3 5...

Task

Time

Predecessor

1

0.2

-

2

0.4

1

3

0.1

1

4

0.4

2,3

5

0.1

2

6

0.3

3,4

7

0.2

5

8

0.2

5,6

9

0.4

7,8

Consider the operations given in Table 1. 47 feet of the oor space along the conveyor is available for installing assembly stations around a closed loop conveyor system. Eeach station is made 6 feet long in order for the operators to work comfortably, and at least 1 feet clearance between stations (except between input and output stations).

  • Find the maximum number of stations on conveyor when the target production rate is 90 units per hr.
  • Find the maximum number of stations on conveyor when the target production rate is 120 units per hr.
  • Find the speed of conveyor when target production rate is 90 units per hr.
  • Find the speed of conveyor when target production rate is 120 units per hr.
  • Find the task to workstation assignments with Largest Candidate Rule when the target production rate is 90 units per hr.
  • Find the total space occupied by the workstations and clearances when the target production rate is 90 units per hr.
  • Find the task to workstation assignments with Largest Candidate Rule when the target production rate is 120 units per hr. Adjust the station-sizes using the method we covered in the class.
  • Find the allocated space occupied to each workstation, and total space occupied including clearances when the target production rate is 120 units per hr.
Answer #1

The Total Task time, T is 2.3 mins, as given.

When the target production rate is 90 units per hr (1.5 units per min), the Cycle time, C (Production Time/ Output) is 0.67 mins.

The theoretical minimum number of stations on conveyor, Nt is T/C rounded up to the next highest integer.

Hence, Nt = 2.3/0.67 = 3.45 = 4 stations

The corresponding Efficiency, Et = T/ (Nt x C) = 0.86

When the target production rate is 120 units per hr (2 units per min), the Cycle time, C is 0.5 mins.

Hence, the theoretical minimum number of stations on conveyor, Nt = 2.3/0.5 = 4.6 = 5 stations

The corresponding Efficiency, Et = 0.92

The task to workstation assignments with Largest Candidate Rule when the target production rate is 90 units per hr, is as tabulated below. Based on this, there are 4 stations (Na) for which the space needed is 27 feet (given that each station is 6 feet long, and there is 1 foot clearance between stations - except between input and output stations.

The corresponding actual Efficiency, E = T/ (Na x C) = 0.86

Task Task time Remaining unassigned time Feasible remaining tasks
WS1 1 0.2 0.466666667 2
2 0.4 0.066666667 None
WS2 3 0.1 0.566666667 4
4 0.4 0.166666667 5
5 0.1 0.066666667 None
WS3 6 0.3 0.366666667 7, 8
7 0.2 0.166666667 None
WS4 8 0.2 0.466666667 9
9 0.4 0.066666667

The task to workstation assignments with Largest Candidate Rule when the target production rate is 120 units per hr, is as tabulated below. Based on this, there are 6 stations for which the space needed is 41 feet (given that each station is 6 feet long, and there is 1 foot clearance between stations - except between input and output stations.

The corresponding actual Efficiency, E = 0.77

Task Task time Remaining unassigned time Feasible remaining tasks
WS1 1 0.2 0.3 3
3 0.1 0.2 None
WS2 2 0.4 0.1 5
5 0.1 0 None
WS3 4 0.4 0.1 None
WS4 6 0.3 0.2 7, 8
7 0.2 0 None
WS5 8 0.2 0.3 None
WS6 9 0.4 0.1